# Reason-to-Recommend: Using Interaction-of-Thought Reasoning to Enhance LLM Recommendation

### v1

@@ 1 Introduction
The integration of Large Language Models (LLMs) into recommendation tasks has recently garnered substantial research interest [1, 2, 3, 4, 5, 6, 7, 8], owing to their strong semantic understanding, contextual reasoning capabilities, and extensive world knowledge acquired through large-scale pretraining. In parallel, enhancing the reasoning capabilities of LLMs has emerged as a key objective [9, 10, 11, 12], with increasing efforts devoted to equipping LLMs with the ability to perform structured, multi-step inference, which is a property that not only improves task performance but also brings enhanced interpretability by making the decision-making process more transparent.
@@ 2.2 LLM reasoning
Recent advances demonstrate that Large Language Models (LLMs) can significantly enhance reasoning capabilities through reinforcement learning (RL)[19, 20, 21, 22] and test-time scaling techniques such as Monte Carlo Tree Search (MCTS)[36, 37, 38, 39] and multi-path sampling [40, 41, 42]. RL-based post-training methods, exemplified by DeepSeek-R1 [24] and Kimi k1.5 [43], facilitate emergent multi-step reasoning behaviors including planning and self-verification, even without supervised initialization. Hybrid approaches like ReFT integrate supervised fine-tuning with RL to further refine reasoning quality. Despite these advances, research on LLM reasoning predominantly targets structured domains such as mathematics and programming, leaving recommendation systems—characterized by implicit user behavior and scarce explicit supervision—largely unexplored, highlighting a critical area for future investigation.

### v2

@@ 1 Introduction
The integration of Large Language Models (LLMs) into recommendation tasks has recently garnered substantial research interest [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], owing to their strong semantic understanding, contextual reasoning capabilities, and extensive world knowledge acquired through large-scale pretraining. In parallel, enhancing the reasoning capabilities of LLMs has emerged as a key objective [15, 16, 17, 18, 19], with increasing efforts devoted to equipping LLMs with the ability to perform structured, multi-step inference, which is a property that not only improves task performance but also brings enhanced interpretability by making the decision-making process more transparent.
@@ 2.2 LLM reasoning
As Large Language Models (LLMs) are increasingly deployed across various domains [43, 44, 45], recent advances demonstrate that they can significantly enhance reasoning capabilities through reinforcement learning (RL)[26, 27, 28, 29, 19] and test-time scaling techniques such as Monte Carlo Tree Search (MCTS)[46, 47, 48, 49] and multi-path sampling [50, 51, 52]. RL-based post-training methods, exemplified by DeepSeek-R1 [31] and Kimi k1.5 [53], facilitate emergent multi-step reasoning behaviors including planning and self-verification, even without supervised initialization. Hybrid approaches like ReFT integrate supervised fine-tuning with RL to further refine reasoning quality. Despite these advances, research on LLM reasoning predominantly targets structured domains such as mathematics and programming, leaving recommendation systems—characterized by implicit user behavior and scarce explicit supervision—largely unexplored, highlighting a critical area for future investigation.